1. Analysis of a sample of a compound of iron and chlorine gas gave 44.06% iron and 55.94% chlorine. Another sample was analyzed to contain 31.52g of iron and 60.03g of chlorine.


a. Is the second sample the same compound as the first sample? show your calculation to support your answer

b. Could the second sample be another compound of iron and chlorine? Support your answer with calculation


1
Expert's answer
2021-09-22T00:26:57-0400

No the second sample is not the same as the first sample

Total percent of first sample "=44.06+55.94=100" %


Total percent of second sample "=31.52+60.03=91.55"%


With chlorine, iron forms another group of industrially important compounds: ferrous chloride, FeCl2; and ferric chloride, FeCl3. Ferrous chloride is obtained as yellow-green deliquescent (moisture-absorbing) crystals by passing dry hydrogen chloride gas over red-hot iron.